OpenStudy (anonymous):

What is the Maclaurin series for (1+x)^(1/x)?

OpenStudy (abb0t):

A maclaurin series is a Taylor series expansion of a function about 0. So this will be long. But try and look for a pattern. Start by taking a few derivatives.

OpenStudy (ghazi):

series expansion at x=0, 1 or what?

OpenStudy (ghazi):

it is expansion about zero \[e-\frac{ ex }{ 2 }+\frac{ 11ex^2 }{ 24 }-\frac{ 7ex^3 }{ 16 }+ ......\]

OpenStudy (abb0t):

\[f(x) = f(0) + f'(0)x + \frac{ f''(0) }{ 2! }x^2+\frac{ f^{(3)}(0) }{ 3! }x^3 \]

OpenStudy (abb0t):

and so forth.

OpenStudy (zehanz):

Maybe the problem is finding the derivatives? Start by writing \[f(x)=(1+x)^{\frac{ 1 }{ x }}\]as\[f(x) = e ^{\ln(1+x)^{\frac{ 1 }{ x }}}=e ^{\frac{ 1 }{ x }\ln(1+x)}\]Now the first derivative is:\[f'(x)=e ^{\frac{ 1 }{ x }\ln(1+x)}(-\frac{ 1 }{ x^2 }\ln(1+x)+\frac{ 1 }{ x }\frac{ 1 }{ 1+x })\]\[=\frac{ 1 }{ x }\left( \frac{ 1 }{ 1+x }-\frac{ 1 }{ x }\ln(1+x) \right)e ^{\frac{ 1 }{ x }\ln(1+x)}\]Hmm, finding f'(0) is not *that* straightforward...